a) 11,254.1 g + 0.1983 g

b) 66.59 L − 3.113 L

c) 8.16 m × 5.1355

d) 0.0154 kg ÷ 88.3 mL

e) 2.64 × 103 cm + 3.27 × 102 cm

work mustbe written


1
Expert's answer
2021-01-18T01:48:47-0500

a) 11,254.1 g + 0.1983 g = 11,254.2983g

b) 66.59 L − 3.113 L = 63.477L

c) 8.16 m × 5.1355 = 41.9057m

d) 0.0154 kg ÷ 88.3 mL = 0.0154kg/mL

e) 2.64 × 103 cm + 3.27 × 102 cm =

271.92cm + 333.54cm = 605.46cm